A 400-g block with an initial speed of 80 cm/s slides along a horizontal tabletop against a friction force of 0.70 N. (a) How far will it slide before stopping? (b) What is the coefficient of friction between the block and the tabletop?


1
Expert's answer
2021-09-01T11:20:00-0400

Given:

"m=400\:\rm g=0.4\: kg"

"v=80\:\rm cm/s=0.8\: m/s"

"F_f=0.70\:\rm N"


(a) The energy-work theorem says

"0-\frac{mv^2}{2}=W=-F_fd"

Hence

"d=\frac{mv^2}{2F}=\frac{0.4*0.8^2}{2*0.7}=0.18\:\rm m"

(b) The friction force

"F_f=\mu N=\mu mg"

Hence

"\mu=\frac{F_f}{mg}=\frac{0.7}{0.4*9.8}=0.18"